Calculate the velocity of an electron if it has a wavelength of 20.8 A O a. 3.6 x 10-25 m/s b. 1.8 x 10-12 m/s c. 3.6x 10+5 m/s O d. 3.6 x 10*6 m/s
Answer #1

\lambda    = 20.8A0

       = 20.8*10^-10m             [ 1A0    = 10^-10 m]

m   = 9*10^-31Kg

h   = 6.626*10^-34 j-sec

\lambda    = h/mv

V   = h/m\lambda

     = 6.626*10^-34 /9*10^-31*20.8*10^-10    =3.6*10^5 m/sec >>>>answer

c. 3.6*10^5 m/sce>>>answer